# Beyond the Label Itself: Latent Labels Enhance Semi-supervised Point Cloud Panoptic Segmentation

### 13 Dec 2023 v1

@@ Introduction
Table 1: Segmentation results compared with other semi-supervised methods on nuScenes and SemanticKITTI. The symbol * in the table denotes semi-supervised results at a ratio of 40%, distinguishing them from other methods at 50%..
@@
Baseline and Implementation Details
Our baseline comprises an image encoder (SwiftNet), a LiDAR encoder (Cylinder3D), and a decoder (Panoptic-PolarNet). It can be comprehended as the combination of section of Image Branch and LiDAR Branch pipeline, as well as Multi-modal Segmentation Network, with the exclusion of the description of the proposed IPSL module and Cylinder-Mix. Additionally, the instance augmentation of PUPS (Su et al. 2023) is applied. More Implementation Details of Detection Network and point cloud segmentation network are in the Appendix.
@@
Firstly, we present the Quantitative Results as table 1. It shows segmentation results on the validation set of nuScenes and SemanticKITTI. As there are limited panoramic segmentation methods available for SS methods, we compare our model with previous methods (French et al. 2019; Tarvainen and Valpola 2017; Zou et al. 2018; Chen et al. 2021; Kong et al. 2023) with semantic segmentation metrics (mIoU, %) and attach our panoptic segmentation metrics (PQ, %). Consequently, our model outperforms the previous top, LaserMix, and attains state-of-the-art semantic results.
@@
Table 3: Ablation of IPSL and Cylinder-Mix on nuScenes
@@
We conduct ablations on the Instance Position-scale Learning (IPSL) Module and Cylinder-Mix (CM) with their default settings. In IPSL Module, the radius of Gaussian heatmaps $R=5$, percentage of scale to the center $P_{center}=1/4$, and the region size and the probability of CM are $[4,4,2]$ and 25%, respectively.
@@ Appendix A Detailed Experiments
Point Cloud Panoptic Segmentation. Following the same configuration of Panoptic-PolarNet, the points among the area of $\{(x,y,z)\mid(3\leq\sqrt{x^{2}+y^{2}}\leq 50)\bigcap(-3\leq z\leq 1.5)\}$ are voxelized to cylinder voxels with the grid size [480, 360, 32]. For Cylinder-Mix, the region size $[R_{x},R_{y},R_{z}]$ is set to [4, 4, 2]. The weights of each loss term in $\mathcal{L}_{seg}$ are as follows: $\mu_{hm}=100,\mu_{os}=10,\mu_{fm}=1$. When training, we set the learning rate at 0.002 and 0.004 for the datasets of SemanticKITTI and nuScenes, respectively. During the pretraining phase, as discussed in the Self-Training section of the paper, we configure 240, 300, 400, and 800 epochs for the experiments of semi-supervised ratios of 40%, 20%, 10%, and 1% on nuScenes. Similarly, for SemanticKITTI, we set 100, 180, 260, and 680 epochs for the corresponding SS ratios. While the epochs of the retraining phase are always 25.
Detection Network of Instance Position-scale Learning Module. The detector is trained on MMdetection (Chen et al. 2019) offline after generating and saving instance boxes in images by 3D-2D projection. We choose Faster R-CNN (Ren et al. 2015) as the backbone and train with a learning rate of 0.002 for SmanticKITTI and 0.005 for nuScenes images. Besides, we set 0.65 as the RPN threshold and 0.45 as the RCNN threshold. Fourth, infer image boxes from trained models, save these boxes, and wait for the fusion of SwiftNet.
@@
Therefore, for a fair and intuitive comparison, we introduce a new evaluation method to precisely assess the evaluation of IPSL on SemanticKITTI, by disregarding the points outside the camera view. Because IPSL primarily affects the points that can be projected onto the camera pixels, we only calculate the PQ metrics for these points. As shown in PQ metrics in Table 2, through limited gain from the IPSL module for all points, it shows a noticeable improvement for points within the camera’s field of view. This validates that IPSL primarily enhances the fusion of 2D-3D features and is actually effective.
@@
Table 4: Ablation of Cylinder-Mix on 1% SemanticKITTI
@@
Firstly, to make a fair comparison with Lasermix, ring-mixing in Lasermix is performed as data augmentation on our baseline instead of Cylinder-Mix. At 25%, LaserMix has a PQ of 46.76%, while Cylinder-Mix can achieve a maximum PQ of 48.97%. Secondly, we fix the region size to [4,4,2] and adjusted the $p_{cylmix}$. The results demonstrate that $p_{cylmix}=1$ is not absolutely optimal and the model tends to benefit more from partially mixed data with original data rather than entirely mixed data. Finally, we fixed the $p_{cylmix}$ to 25% and adjusted the region size to [4,4,2], [16,16,4], [1,1,4] and [8,8,1], respectively. It is expected to evaluate the sensitivity of Cylinder-Mix along with the Z-axis and X-Y axes. We don’t separately increase the region size of the X and Y axes because they are equivalent in the cylinder voxel. Based on the experimental results, we draw the conclusion that adding region size solely at $R_{z}$ or the XY-plane is not the optimal mixing strategy, and too few regions to mix are also not ideal. A moderately large region size, such as [4,4,2], is a good choice of region size.
@@ Appendix C Visualization
As shown in Figure 2, we provide visualizations of point clouds for the ground truth, predictions of the baseline, and our model. The color red is used to indicate incorrect predictions in the point cloud. Semantic segmentation and instance segmentation visualisations are presented. The mIoU or PQ excluding categories that are absent within ground truth is labeled in the bottom-right corner of the subplot. The 100% on the left indicates the model trained on the full SemanticKITTI dataset, while the 10% on the right means only using on 10% labeled subset following our semi-supervised setting.
Table 6: Results on Detection Network of different backbone
@@
Table 7: Class-wise IoU and PQ scores on the val set of SemanticKITTI.

### 11 Feb 2024 v2

@@ Introduction
Table 1: Segmentation results compared with other semi-supervised methods on validation set of nuScenes and SemanticKITTI. The symbol * in the table denotes semi-supervised results at a ratio of 40%, distinguishing them from other methods at 50%.
@@
Implementation Details
Our baseline comprises an image encoder (SwiftNet), a LiDAR encoder (Cylinder3D), and a decoder (Panoptic-PolarNet). It can be comprehended as the combination of section of Image Branch and LiDAR Branch pipeline, as well as Multi-modal Segmentation Network, with the exclusion of the description of the proposed IPSL module and Cylinder-Mix. Additionally, the instance augmentation of PUPS (Su et al. 2023) is applied.
Implementation Default settings unless ablation. In IPSL Module, the radius of Gaussian heatmaps $R=5$, percentage of scale to the center $P_{center}=1/4$. For Cylinder-Mix, the region size $[R_{x},R_{y},R_{z}]=[4,4,2]$, the probability $p_{clymix}=25\%$. For the weight of each term in loss in $\mathcal{L}_{seg}$, $\mu_{hm}=100,\mu_{os}=10,\mu_{fm}=1$. More Implementation Details of Detection Network and point cloud segmentation network are in the Appendix.
@@
Firstly, we present the Quantitative Results as Table 1. It shows segmentation results on the validation set. As there are limited panoramic segmentation methods available for SS methods, we compare our model with previous methods (French et al. 2019; Tarvainen and Valpola 2017; Zou et al. 2018; Chen et al. 2021; Kong et al. 2023) with semantic segmentation metrics (mIoU, %) and attach our panoptic segmentation metrics (PQ, %). Eventually, our model outperforms the previous top, LaserMix, and attains state-of-the-art semantic results.
@@
Table 2: Comparison of Cylinder-Mix between LaserMix.
@@
We conduct ablations on the Instance Position-scale Learning (IPSL) Module and Cylinder-Mix (CM) with their default settings.
@@ Appendix A Detailed Experiments
Point Cloud Panoptic Segmentation. Following the same configuration of Panoptic-PolarNet, the points among the area of $\{(x,y,z)\mid(3\leq\sqrt{x^{2}+y^{2}}\leq 50)\bigcap(-3\leq z\leq 1.5)\}$ are voxelized with the grid size [480, 360, 32]. For Cylinder-Mix, the region size $[R_{x},R_{y},R_{z}]=[4,4,2]$. the weights of each term in loss $\mathcal{L}_{seg}$, $\mu_{hm}=100,\mu_{os}=10,\mu_{fm}=1$. Besides, we set the learning rate at 0.002 and 0.004 for the datasets of SemanticKITTI and nuScenes, respectively. During the pretraining phase, as discussed in the Self-Training section of the paper, we configure 240, 300, 400, and 800 epochs for the experiments of semi-supervised ratios of 40%, 20%, 10%, and 1% on nuScenes. Similarly, for SemanticKITTI, we set 100, 180, 260, and 680 epochs for the corresponding SS ratios. While the epochs of the retraining phase are always 25.
Detection Network. The detector is trained on MMdetection (Chen et al. 2019) offline after generating and saving instance boxes in images by 3D-2D projection. We choose Faster R-CNN (Ren et al. 2015) as the backbone and train with a learning rate of 0.002 for SmanticKITTI and 0.005 for nuScenes images. Besides, we set 0.65 as the RPN threshold and 0.45 as the RCNN threshold. Fourth, infer image boxes from trained models, save these boxes, and wait for the fusion of SwiftNet.
Module Cost On nuScenes, the number parameters of baseline is 296.2MB, and the memory cost is 26.2GB (batch size=2). With Cylinder-Mix, the memory and time cost (1.2 seconds per frame) remain nearly unchanged. When the IPSL module is added, memory and time costs acceptably increase by 3.5GB and 0.13 seconds per frame, respectively.
@@
Therefore, we introduce a new evaluation method for a fair comparison. To precisely assess the evaluation of IPSL on SemanticKITTI, we only calculate the PQ metrics for these points inside the camera view, cause IPSL primarily affects the points that can be projected onto the camera pixels. As shown in PQ metrics in Table 2, through limited gain from the IPSL module for all points, it shows a noticeable improvement for points within the camera’s field of view. This validates that IPSL primarily enhances the fusion of 2D-3D features and is actually effective.
@@
Table 3: Ablation study on boxes heatmap
@@
Firstly, to make a fair comparison with Lasermix, ring-mixing in Lasermix is performed as data augmentation on our baseline instead of Cylinder-Mix. At 25%, LaserMix has a PQ of 47.0%, while CM can achieve a maximum PQ of 49.0%. Secondly, we fix the region size to [4,4,2] and adjusted the $p_{cylmix}$. The results demonstrate that $p_{cylmix}=1$ is not absolutely optimal and the model tends to benefit more from partially mixed data with original data rather than entirely mixed data. Finally, we fixed the $p_{cylmix}$ to 25% and adjusted the region size to [4,4,2], [16,16,4], [1,1,4] and [8,8,1], respectively. It is expected to evaluate the sensitivity of Cylinder-Mix along with the Z-axis and X-Y axes. We don’t separately increase the region size of the X and Y axes because they are equivalent in the cylinder voxel. Based on the experimental results, we draw the conclusion that adding region size solely at $R_{z}$ or the XY-plane is not the optimal mixing strategy, and too few regions to mix are also not ideal. A moderately large region size, such as [4,4,2], is a good choice of region size.
@@ Appendix C Visualization
As shown in Figure 2, we provide visualizations of point clouds for the ground truth, predictions of the baseline, and our model. The color red is used to indicate incorrect predictions in the point cloud. Semantic segmentation and instance segmentation visualisations are presented. The mIoU or PQ excluding categories that are absent within ground truth is labeled in the bottom-right corner of the subplot. The 100% on the left indicates the model trained on the full SemanticKITTI dataset, while the 10% on the right means only using on 10% labeled subset following our SS setting.
@@
Table 6: Results on Detection Network of different backbone